1. A soap bubbles 50mm diameter has internal pressure in excess of the outside pressure of 25N/m. Calculate the tension in soap film

2. Determine the minimum size of glass tubing that can be used to measure water level. If the capillary rise in the tube does not exceed 0.25mm. Take surface tension of water in contact with air as 0.0735N/m
1
Expert's answer
2020-05-24T18:06:10-0400

1)


"p=\frac{8\sigma}{d}\to 25=\frac{8\sigma}{0.05}"

"\sigma=0.156\frac{N}{m}"

2)

"h=\frac{4\sigma}{\rho gd}\to 0.25\cdot10^{-3}=\frac{4(0.0735)}{(1000)(9.81)d}"

"d=0.120\ m=12.0\ cm"